a) Determine whether or not the following are subspaces?

i. 𝑾 = {(𝒂,𝒃, 𝒄) ∈ β„πŸ‘

|𝒂 + 𝒃 + 𝒄 = 𝟎} of β„πŸ‘

ii. The symmetric matrices of 𝑴𝒏𝒏 (the vector space of 𝒏 Γ— 𝒏

matrices)

iii. All polynomials of degree 2.

b)Β 

i. For which real values of 𝝀 do the following vectors form aΒ 

linearly dependent set in β„πŸ‘

?

π’—πŸ = (𝝀, βˆ’

𝟏

𝟐

, βˆ’

𝟏

𝟐

) , π’—πŸ = (βˆ’

𝟏

𝟐

, 𝝀, βˆ’

𝟏

𝟐

) , π’—πŸ‘ = (βˆ’

𝟏

𝟐

, βˆ’

𝟏

𝟐

, 𝝀)

ii. Find a basis and dimension of the solution space for theΒ 

following homogenous linear equations:Β 

π’™πŸ + πŸπ’™πŸ βˆ’ π’™πŸ‘ + πŸ’π’™πŸ’ = 𝟎

πŸπ’™πŸ βˆ’ π’™πŸ + πŸ‘π’™πŸ‘ + πŸ‘π’™πŸ’ = 𝟎

πŸ’π’™πŸ + π’™πŸ + πŸ‘π’™πŸ‘ + πŸ—π’™πŸ’ = 𝟎

π’™πŸ βˆ’ π’™πŸ‘ + π’™πŸ’ = 𝟎

πŸπ’™πŸ + πŸ‘π’™πŸ βˆ’ π’™πŸ‘ + πŸ•π’™πŸ’ = 𝟎



1
Expert's answer
2021-06-22T02:42:24-0400

a)

i.

"W=\{(a,b,c)\in R^3|a+b+c=0\} of R^3\\
x=(a,b,c), a+b+c=0\\
y=(a_1,b_1,c_1), a_1+b_1+c_1=0\\
(1)x+y=(a+a_1,b+b_1,c+c_1), \\
a+a_1+b+b_1+c+c_1=0+0=0\\
x+y\in W\\
i) x+y=y+x\\
ii) (x+y)+z=x+(y+z)\\
iii) \exists 0=(0,0,0)\in W, x+0=x\\
iiii) \forall x(a,b,c)\in W \\
\exists (-x)=(-a,-b,-c)\in W (-a+(-b)+(-c)=-(a+b+c)=0):\\
x+(-x)=0"

"(2) r\in R, x=(a,b,c)\in W, a+b+c=0\\
r\cdot x=r(a,b,c)=(ra,rb,rc)\in W\\
ra+rb+rc=r(a+b+c)=0\\
i) r,s\in R: (r+s)x=r\cdot x+s\cdot x\\
ii) r(x+y)=r\cdot x+r\cdot y\\
iii) (rs)x=r \cdot(s\cdot x)\\
iiii)1\cdot x=x"

W are subspaces


ii.

"M_{nn}=\{\begin{pmatrix}
 a_{11} & a_{12}&...&a_{1n} \\
 a_{12} & a_{22}&...&a_{2n}\\
...&...&...&...\\
a_{1n} & a_{2n}&...&a_{nn} 
\end{pmatrix}, a_{ij}\in R\}\\
A=\begin{pmatrix}
 a_{11} & a_{12}&...&a_{1n} \\
 a_{12} & a_{22}&...&a_{2n}\\
...&...&...&...\\
a_{1n} & a_{2n}&...&a_{nn} 
\end{pmatrix}\in M_{nn}\\
B=\begin{pmatrix}
 b_{11} & b_{12}&...&b_{1n} \\
 b_{12} & b_{22}&...&b_{2n}\\
...&...&...&...\\
b_{1n} & b_{2n}&...&b_{nn} 
\end{pmatrix}\in M_{nn}"

"A+B=\begin{pmatrix}
 a_{11} & a_{12}&...&a_{1n} \\
 a_{12} & a_{22}&...&a_{2n}\\
...&...&...&...\\
a_{1n} & a_{2n}&...&a_{nn} 
\end{pmatrix}+\\+\begin{pmatrix}
 b_{11} & b_{12}&...&b_{1n} \\
 b_{12} & b_{22}&...&b_{2n}\\
...&...&...&...\\
b_{1n} & b_{2n}&...&b_{nn} 
\end{pmatrix}=\\
=\begin{pmatrix}
 a_{11}+b_{11} & a_{12}+b_{12}&...&a_{1n}+b_{1n} \\
 a_{12} + b_{12}& a_{22}+ b_{22}&...&a_{2n}+ b_{2n}\\
...&...&...&...\\
a_{1n}+ b_{1n} & a_{2n}+ b_{2n}&...&a_{nn}+b_{nn} 
\end{pmatrix}\in M_{nn}"

"i) A+B=B+A\\
ii) (A+B)+C=A+(B+C)\\
iii) \exists 0=\begin{pmatrix}
 0 & 0&...&0 \\
 0 & 0&...&0\\
...&...&...&...\\
 0 & 0&...&0
\end{pmatrix}\in W_{nn}, A+0=A\\"


"iiii) \forall A=\begin{pmatrix}
 a_{11} & a_{12}&...&a_{1n} \\
 a_{12} & a_{22}&...&a_{2n}\\
...&...&...&...\\
a_{1n} & a_{2n}&...&a_{nn} 
\end{pmatrix}\in M_{nn},\\
\exists (-A)=\begin{pmatrix}
 -a_{11} & -a_{12}&...&-a_{1n} \\
 -a_{12} & -a_{22}&...&-a_{2n}\\
...&...&...&...\\
-a_{1n} & -a_{2n}&...&-a_{nn} 
\end{pmatrix} \in M_{nn}:\\
A+(-A)=0"

"(2) r\in R, A\in M_{nn}\\
r\cdot A=r\cdot \begin{pmatrix}
 a_{11} & a_{12}&...&a_{1n} \\
 a_{12} & a_{22}&...&a_{2n}\\
...&...&...&...\\
a_{1n} & a_{2n}&...&a_{nn} 
\end{pmatrix}=\\
=\begin{pmatrix}
 r\cdot a_{11} &r\cdot a_{12}&...&r\cdot a_{1n} \\
 r\cdot a_{12} & r\cdot a_{22}&...&r\cdot a_{2n}\\
...&...&...&...\\
r\cdot a_{1n} & r\cdot a_{2n}&...&r\cdot a_{nn} 
\end{pmatrix}\in M_{nn}"


"i) r,s\in R: (r+s)A=r\cdot A+s\cdot A\\
ii) r(A+B)=r\cdot A+r\cdot B\\
iii) (rs)A=r \cdot(s\cdot A)\\
iiii)1\cdot A=A"

"M_{nn}" are subspaces


iii.

"P_2=\{ax^2+bx+c| a,b,c\in R\}\\
f=ax^2+bx+c\in P_2\\
f_1=a_1x^2+b_1x+c_1\in P_2\\
(1)f+f_1=ax^2+bx+c+a_1x^2+b_1x+c_1=\\
=(a+a_1)x^2+(b+b_1)x+(c+c_1)\in P_2"

"i) f+f_1=f_1+f\\
ii) (f+f_1)+f_2=f+(f_1+f_2)\\ 
iii) \exists 0=0\cdot x^2+0\cdot x+0)\in P_2, f+0=f\\
iiii) \forall f\in P_2 \\
\exists (-f)=-ax^2+(-b)x+(-c)\in P_2:\\
f+(-f)=0"

"(2) r\in R, f=ax^2+bx+c\in P_2\\
r\cdot f=r(ax^2+bx+c)=\\
=r\cdot ax^2+r\cdot bx+r\cdot c\in P_2\\
i) r,s\in R: (r+s)f=r\cdot f+s\cdot f\\
ii) r(f+f_1)=r\cdot f+r\cdot f_1\\
iii) (rs)f=r \cdot(s\cdot f)\\
iiii)1\cdot f=f"

"P_2" are subspaces


b)

i.

"v_1=(\lambda,-12,-12)\\
v_2= (-12,\lambda,-12)\\
v_3=(-12,-12,\lambda)\\
a\cdot v_1+b\cdot v_2+c\cdot v_3=0\\
a\cdot \lambda-12b-12c=0\\
-12a+b\cdot \lambda-12c=0\\
-12a-12b+c\cdot \lambda=0"

"\Delta=\begin{vmatrix}
 \lambda & -12&-12 \\
 -12&\lambda& -12\\
-12&-12&\lambda 
\end{vmatrix}=\lambda ^3-432\lambda -3456\neq0\\
\lambda ^3-432\lambda -3456=0\\
\lambda ^3+12\lambda^2-12\lambda^2-432\lambda -3456=0\\
\lambda ^2(\lambda+12)-12(\lambda^2+36\lambda +288)=0\\
\lambda ^2(\lambda+12)-12(\lambda+12)(\lambda +24)=0\\
(\lambda+12)(\lambda ^2-12\lambda-288)=0\\
\lambda=-12, \lambda=-12,\lambda=24"

For "\lambda=-12, \lambda=24" the following vectors form aΒ 

linearly dependent set in "R^3".


ii.

"x_1+2x_2-x_3+4x_4=0\\
2x_1-x_2+3x_3+3x_4=0\\
4x_1+x_2+3x_3+9x_4=0\\
x_2-x_3+x_4=0\\
2x_1+3x_2-x_3+7x_4=0\\
\begin{pmatrix}
 1& 2&-1&4&|0 \\
 2&-1&3&3&|0\\
4&1&3&9&|0\\
0&1&-1&1&|0\\
2&3&-1&7&|0
\end{pmatrix}\\
IIr+Ir(-2)\\IIIr+Ir(-4)\\IIIIr+Ir(-2)\\"

"\begin{pmatrix}
 1& 2&-1&4&|0 \\
 0&-5&5&-5&|0\\
0&-7&7&-7&|0\\
0&1&-1&1&|0\\
0&-3&3&-6&|0
\end{pmatrix}\\
III+IIIIr\cdot 5\\IIIr+IIIIr\cdot 7\\IIIIr+IIIIr \cdot 3"

"\begin{pmatrix}
 1& 2&-1&4&|0 \\
 0&1&-1&1&|0\\
0&0&0&0&|0\\
0&0&0&0&|0
\end{pmatrix}\\"

"x_1+2x_2-x_3+4x_4=0\\
x_2-x_3+x_4=0\\
x_1=-x_3-2x_4\\
x_2=x_3-x_4"

Basis

"x_1=(-1,1,1,0)\\
x_2=(-2,-1,0,1)"

dim =2